Define the variables
State and justify the assumption
Explain with words flow of logic for each step in the problem
.Il Verizon令 오후 4:21 Back Problem 2 Option C Application Problem 2, Option C An oil tankers engines have broken down, and the wind is blowing the tanker straight toward a reef at a constant speed of 1.5 m/s. When the tanker is 500 m from the reef, the wind dies down just as the engineer gets the engines going again. The rudder is stuck, so the only choice is to try to accelerate straight backward away from the reef. The mass of the tanker and cargo is 3.6x10 kg, and the engines produce a net horizontal force of 8.0x10 N on the tanker a.) Will the ship hit the reef? b.) If it does, will the oil be safe? The hull can withstand an impact speed of 0.2 m/s or less. You can ignore the retarding force of the water on the tankers hull. Previous Next Dashboard Calendar To Do Notifications Inbox
Answer #1

a] Acceleration provided by the engine = force provided by the tanker / mass of tanker and cargo

a = F/m = (8 x 104)/(3.6 x 107) = 0.00222 m/s2

this acceleration is to the left (assuming the tanker is initially moving to the right)

the force provided here will first momentarily bring the tanker to rest and then accelerate it to the left.

use v2 = u2 + 2aS

v2 = (1.5)2 - 2(0.00222)(500) = 0.02777

=> v = 0.1667 m/s

since this velocity is not zero, the ship will hit the reef.

b] The hull can withstand an impact speed of 0.2 m/s. Since the impact speed here is 0.1667 m/s, the oil will be safe.